### v1

# Bayesian spatio-temporal weighted regression for integrating missing and misaligned environmental data

Abstract
Estimating environmental exposures from multi-source data is central to public health research and policy. Integrating data from satellite products and ground monitors are increasingly used to produce exposure surfaces. However, spatio-temporal misalignment often induced from missing data introduces substantial uncertainty and reduces predictive accuracy. We propose a Bayesian weighted predictor regression framework that models spatio-temporal relationships when predictors are observed on irregular supports or have substantial missing data, and are not concurrent with the outcome. The key feature of our model is a spatio-temporal kernel that aggregates the predictor over local space-time neighborhoods, built directly into the likelihood, eliminating any separate gap-filling or forced data alignment stage. We introduce a numerical approximation using a Voronoi-based spatial quadrature combined with irregular temporal increments for estimation under data missingness and misalignment. We showed that misspecification of the spatial and temporal lags induced bias in the mean and parameter estimates, indicating the need for principled parameter selection. Simulation studies confirmed these theoretical findings, where careful tuning was critical to control bias and achieve accurate prediction, while the proposed quadrature performed well under severe missingness. As an illustrative application, we estimated fine particulate matter (PM2.5) in northern California using satellite-derived aerosol optical depth (AOD) and wildfire smoke plume indicators. Relative to a traditional collocated linear model, our approach improved out-of-sample predictive performance (over 50% increase in R2), reduced uncertainty, and yielded robust temporal predictions and spatial surface estimation. Our framework is extensible to additional spatio-temporally varying covariates and other kernel families.
Keywords Spatio-temporal regression $\cdot$ Kernel-averaged predictor $\cdot$ Missing and misaligned data $\cdot$ PM2.5 mapping
1 Introduction
Environmental exposures derived from multi-source data play a central role in public health. Accurate exposure estimates are needed to quantify population-level risks, guide interventions, and inform regulatory action [Clark et al., 2025]. Data streams such as satellite observations, ground monitors, and physical model outputs provide complementary information, but they often differ in spatial resolution, temporal frequency, and coverage [Lolli, 2025]. Integrating these sources is essential for many applications but can be methodologically challenging [Forlani et al., 2020]. Missing observations, spatial or temporal misalignment, and nonstationary associations between predictors and outcomes introduce uncertainty, limiting predictive reliability and the robustness of scientific inference.
Air quality research provides a clear example of these challenges. Exposure to fine particulate matter (PM2.5) is a leading risk factor for respiratory and cardiovascular disease and premature mortality worldwide, contributing to an estimated 8.1 million deaths globally in 2021 [Health Effects Institute, 2024]. While ground-based monitoring networks provide reliable measurements, they are spatially sparse and unevenly distributed [Zhang et al., 2021], limiting their ability to capture fine-scale spatial variability. Satellite-derived aerosol optical depth (AOD), available from instruments like the Moderate resolution Imaging Spectroradiometer (MODIS), offer global spatial coverage and have been widely used as proxies for surface PM2.5 [Diao et al., 2019]. However, AOD retrievals are often missing due to cloud cover, snow (bright surfaces), or algorithmic failures, and when available, they are not always co-located or spatially aligned with ground monitors, leading to incomplete or misaligned data.
To address these limitations, statistical approaches have been developed to model the PM2.5–AOD relationship and integrate additional sources of data such as meteorology and land use characteristics. Early methods relied on linear regression [Wang and Christopher, 2003, Chu et al., 2003] and were later extended through multiple linear regression and generalized additive models to incorporate meteorological covariates [van Donkelaar et al., 2006, Liu et al., 2009, Franklin et al., 2017]. In these studies, spatial misalignment was commonly addressed by interpolating or averaging AOD onto regular grids first, then modeling the PM-AOD relationship. Machine learning and deep learning have come to the forefront in interpolation efforts as well, but explicit spatio-temporal trends are not generally incorporated in these approaches [Li et al., 2020]. Multi-stage spatio-temporal models have been proposed to estimate daily PM2.5 using neighbouring observations where AOD is missing, [Pu and Yoo, 2020]; however, these approaches remain limited in their ability to capture nonstationarity and locally varying relationships.
Local regression techniques provide one pathway to address this gap. Geographically weighted regression (GWR) [Brunsdon et al., 1998] and its spatio-temporal extension, GTWR [Fotheringham et al., 2015] apply location- and time-specific weights, allowing regression coefficients to vary across both space and time. GTWR has been shown to improve predictive performance for PM2.5 mapping by capturing nonstationary AOD–PM2.5 relationships [He and Huang, 2018, Hu et al., 2013], producing fine-resolution exposure surfaces that reveal gradients and hotspots. Yet, these models still rely on concurrent predictor–outcome pairs, making them vulnerable to the systematic missingness and misalignment common in environmental datasets.
Other methodological advances in spatial modeling have aimed to extend beyond strictly concurrent data. Kernel-averaged predictor frameworks [Heaton and Gelfand, 2011, Heaton and Gelfand, 2012] smoothed predictors over a spatial or spatio-temporal buffer, allowing nearby information to inform estimates even when direct matches are unavailable. While promising, these methods can be sensitive to the choice of kernel size and lag parameters, and their robustness to missing data remains unclear. This is a pressing concern in many applications, where monitoring networks are sparse, satellite retrievals are often intermittent, and episodic events such as wildfire smoke plumes create rapidly evolving exposure patterns.
In this paper, we build on the developments of [Heaton and Gelfand, 2012] by proposing a new method of estimation for the weighted predictor regression framework, tailored to settings with missing or misaligned data. Our Bayesian formulation allows flexible spatial and temporal lag structures, enabling regression even when predictors are not observed concurrently with the outcome. The framework accommodates varying buffer sizes and temporal lag specifications, and we evaluate its robustness to model misspecification as well as its performance under substantial missingness in the predictors. We applied the approach to estimate PM2.5 in Northern California from satellite AOD and wildfire smoke plumes, quantifying local associations and demonstrating gains in predictive performance relative to traditional concurrent data models. The paper is organized as follows: Section 2 describes the methodology, provides theoretical bounds for bias induced by model misspecification, and outlines model fitting procedures. Section 3 presents a simulation study evaluating model sensitivity and robustness. Section 4 applies the methods to study spatio-temporal dependencies between daily PM2.5 and AOD during a year where there was strong influence of wildfires on air quality in the region.
2 Method
We denote a spatio-temporal process by $\{Y(\mathbf{s},t):\mathbf{s}\in\mathbb{R}^{d},t\in\mathbb{R}\}$ where $Y(\mathbf{s},t)$ represents an observation at location $\mathbf{s}$ and time $t$ and assume that $Y(\mathbf{s},t)$ depends on another spatio-temporal temporal process $\{X(\mathbf{u},\tau):\mathbf{u}\in\mathbb{R}^{d},\tau\in\mathbb{R}\}$. In a general setting, the proposed model, similar to [Heaton and Gelfand, 2012], is given by
$\mathrm{Y}(\mathbf{s},t)=\beta_{0}(\mathbf{s},t)+\beta_{1}(\mathbf{s},t)\mathrm{W}(\mathbf{s},t)+\varepsilon(\mathbf{s},t).$ (1)
where the weighted predictor, $\mathrm{W}(\mathbf{s},t)$, is
$\mathrm{W}(\mathbf{s},t)=\int_{0}^{q}\int_{\mathcal{B}_{r}(\mathbf{s})}\kappa(\mathbf{s}-\mathbf{u},t-l\;|\;\xi(\mathbf{s},t))\mathrm{X}(\mathbf{u},t-l)d\mathbf{u}dl$
[…]
This framework captures spatio-temporal variation by accounting for the local weighting of nearby predictors and recent time points, as well as the regression effect. These two sets of parameters play different roles: the (kernel) decay parameters $\xi(\mathbf{s},t)$ determine the spatial extent over which the predictors can influence the parameters, whereas the regression parameters $\beta_{j}(\mathbf{s},t)$ govern the strength of the association between the weighted predictor and the response. We model both sources of non-stationarity using smooth functions or covariates observed at the corresponding locations and times:
$\beta_{j}(\mathbf{s},t)=\sum b_{jk}f_{k}(s,t)\textrm{ and }\mathbb{E}(\xi(\mathbf{s},t))=\sum_{g=1}^{G}e_{g}v_{g}(s,t),\;\xi(\mathbf{s},t)\sim\mathcal{G}(\cdot)$
where $f_{k}$ and $v_{g}$ represent smooth basis functions (e.g. splines), or spatio-temporal covariates. When positivity is required for the kernel parameters, we model $\log(\xi(\mathbf{s},t))$ so that $\xi(\mathbf{s},t)\geq 0$.
2.1 Choice of radius and time lag
The choice of $r$ and $q$ must be treated as a model selection problem in which values that maximize goodness-of-fit measures such as R2 and/or minimize Akaike Information Criteria (AIC) and mean-squared error (MSE) are selected. [Heaton and Gelfand, 2011] choose $r$ based on the effective spatial range of the predictor. However, the effect of misspecifying $r$ and $q$ on parameter estimation and predictive performance remains unexplored. In this work, we focus on kernel functions of the form
[…]
Proposition 1.
Let $X(\mathbf{u},\tau)\geq 0$ be a spatio–temporal process and let
[…]
where $\mathcal{B}_{r}(\mathbf{s})=\{\mathbf{u}:\|\mathbf{s}-\mathbf{u}\|\leq r\}$ and $\kappa_{1},\kappa_{2}$ are nonnegative kernels. Define the true mean
$\mu(\mathbf{s},t)=\beta_{0}+\beta_{1}\,W(\mathbf{s},t;q,r,\boldsymbol{\xi})$
and the misspecified mean
$\tilde{\mu}(\mathbf{s},t)=\beta_{0}+\beta_{1}\,W(\mathbf{s},t;\tilde{q},\tilde{r},\boldsymbol{\xi}).$
Then the difference between the two mean functions satisfies
[…]
where
[…]
is the spatial annulus between the two radii.
Proposition 1 implies that the discrepancy between the true and misspecified means is influenced not only by the behaviour of the kernel functions but also by the magnitude of the process outside the true spatial–temporal window.
Corollary 1.
Suppose there exist constants $\delta_{1},\delta_{2}>0$ such that
[…]
Then, for any $\tilde{r}\geq r$ and $\tilde{q}\geq q$, the difference between the true and misspecified means satisfies
$\bigl|\mu(\mathbf{s},t)-\tilde{\mu}(\mathbf{s},t)\bigr|$ $\leq\beta_{1}\left(\delta_{1}\delta_{2}\int_{0}^{\tilde{q}}\int_{\mathcal{A}_{r,\tilde{r}}(\mathbf{s})}X(\mathbf{u},t-\ell)\,d\mathbf{u}\,d\ell\;+\;\delta_{1}\delta_{2}\int_{q}^{\tilde{q}}\int_{\mathcal{B}_{\tilde{r}}(\mathbf{s})}X(\mathbf{u},t-\ell)\,d\mathbf{u}\,d\ell\right)$
$\leq\beta_{1}\bigg(\delta_{1}\delta_{2}\,\tilde{q}\,\pi(\tilde{r}^{2}-r^{2})\,\sup_{(\mathbf{u},\ell)\in\mathcal{A}_{r,\tilde{r}}(\mathbf{s})\times[0,q]}X(\mathbf{u},t-\ell)\;+\;$
$\delta_{1}\delta_{2}\,(\tilde{q}-q)\,\pi\tilde{r}^{2}\,\sup_{(\mathbf{u},\ell)\in\mathcal{B}_{\tilde{r}}(\mathbf{s})\times[q,\tilde{q}]}X(\mathbf{u},t-\ell)\bigg).$
From collorary 1, if $\delta_{1},\delta_{2}\to 0$, then the bound only depends on the behaviour of $X(\mathbf{u},t-l)$ i.e. as long as $X(\mathbf{u},t-l)$ does not contain extremes, the error bound $\to 0$. In this setting, one might consider selecting $\tilde{r}$ and $\tilde{q}$ such that $\mathcal{B}(\tilde{r})\times[0,\tilde{q}]=\mathcal{D}\times\mathcal{T}$, thereby covering the entire spatio-temporal domain under consideration. While this choice appears reasonable in theory, it is often computationally prohibitive in practice. For example, on a regular 1 km² daily grid, the number of data points required to cover a region of radius $\mathcal{B}(\tilde{r})\times[0,\tilde{q}$] is $\pi\tilde{r}^{2}\tilde{q}$, which grows quadratically with $\tilde{r}$. Thus, it is recommended that model selection be based on performance assessed under a set of varying spatial and temporal lag values.
2.2 Model estimation
2.2.1 Numerical approximation of spatio-temporal integral
In practice, we only observe the spatio-temporal processes at discrete time-points and locations: $\{Y(s_{n},t_{i})\;:i\in\{1,\ldots,\mathrm{T_{Y_{n}}}\},n\in\{1,\ldots,\mathrm{N_{Y}}\}\}$, and $\{X(s_{m},t_{j})\;:j\in\{1,\ldots,\mathrm{T_{X_{m}}}\},m\in\{1,\ldots,\mathrm{N_{X}}\}\}$.
[…]
$Y(\mathbf{s}_{n},t_{i})=\beta_{0}(\mathbf{s}_{n},t_{i})+\beta_{1}(\mathbf{s}_{n},t_{i})\mathrm{W}(\mathbf{s}_{n},t_{i}\;;\;\mathbf{\xi}(\mathbf{s}_{n},t_{i}))+\varepsilon(\mathbf{s}_{n},t_{i}).$ (3)
When data are missing or misaligned, this estimation is more challenging. [Heaton and Gelfand, 2011] use Monte Carlo integration and [Heaton and Gelfand, 2012] project $X(\mathbf{s},t)$ onto a regular grid and approximate the integral by summing pointwise products of the discretised covariate with its weights after fitting a latent Gaussian process - an approach that can be computationally expensive for finely spaced data due to covariance estimation. We instead use a numerical approximation for the weighted predictor function:;
$\mathrm{W}(\mathbf{s}_{n},t_{i})=\sum_{k}^{M_{ni}}\left(\kappa(d_{ni,k}\;|\;\mathbf{\xi}(\mathbf{s}_{n},t_{i}))\cdot\tilde{\mathbf{X}}_{ni,k}\right)\cdot\Delta\tilde{\mathbf{X}}_{ni,k}$
where $\Delta\tilde{\mathbf{X}}_{ni,k}=\textrm{area}(V_{ni,k})\cdot\delta_{k}$. Denote the $k^{th}$ observation in the buffer by $\tilde{\mathbf{X}}_{ni,k}=X(\mathbf{s}_{k},t_{k})$, then $V_{ni,k}$ is the corresponding area of the cell from the Voronoi tessellation of the spatial location $\mathbf{s}_{k}$. Suppose $X(s_{k},\cdot)$ is observed at time-points $(t_{1},\ldots,t_{k-1},t_{k},\ldots,T_{X_{k}})$ then $\delta_{k}=t_{k}-t_{k-1}$.
A Voronoi tessellation is a computational technique that partitions a metric space based on proximity to a given set of points [Okabe et al., 2009]. At time $i$, given a finite set of distinct points given by $\mathbf{z}_{ni}$ which represents the locations at which we have observed data within the buffer of location $n$, the Voronoi cell $V_{ni,k}$ associated with location $z_{ni,k}$ is defined as:
$V_{ni,k}=\left\{x\in\mathcal{B}_{r}(\mathbf{s}_{n})\mid\|x-z_{ni,k}\|\leq\|x-z_{ni,g}\|,\ \forall g\neq k\right\}.$
Each cell $V_{ni,k}$ consists of all points in the plane that are closer to $z_{ni,k}$ than to any other site. The union of all such cells forms a partition of the buffer region. The area of each Voronoi cell $V_{ni,k}$ quantifies the spatial influence or weight of the corresponding observation point, based on proximity to neighbouring points.
Because $Y$ and $X$ are typically collected on different space-time supports, the choice of $r$ and $q$ directly affects $M_{ni}$, the number of covariate values contributing to $W(\mathbf{s}_{n},t_{i})$. In applications such as PM2.5 estimation from satellite AOD, there can be substantial missing data in $X(\mathbf{v},\tau)$ that yield $M_{ni}=0$ (or very small), in which case $Y(\mathbf{s}_{n},t_{i})$ is poorly approximated. In this case, $Y(\mathbf{s}_{n},t_{i})$ should be removed to avoid large approximation error. In Appendix A Figure 7, we illustrate the Voronoi tessellation of the buffer on two different days. The tessellation depends on the location of the observed AOD data.
2.2.2 Bayesian inference
A Metropolis-within-Gibbs sampler is used to perform inference. We assume a Gaussian for prior $\mathbf{\beta}\sim\mathcal{N}(0,\sigma^{2}_{b})$. This choice of prior is conditionally conjugate under a Gaussian process assumption for $\varepsilon(\mathbf{s},t)\sim\mathcal{GP}(0,\sigma^{2}C(\mathbf{s},\mathbf{s}^{\prime},t,t^{\prime}))$. A conditionally conjugate prior $\mathcal{IG}(a_{b},b_{b})$ is assumed for $\sigma_{b}^{2}$. Similarly, a conditionally conjugate non-informative inverse-Gamma prior is assumed for $\sigma^{2}\sim\mathcal{IG}(a,b)$. The prior for $\mathbf{\xi}(\mathbf{s},t)$ depends on the choice of covariance function and assumed model. In general for $\mathbf{\xi}(\mathbf{s},t)=\xi$ , we propose an informative Gamma $\sim\mathcal{G}(c_{1},c_{2})$ for non-spatio-temporal scale parameters governing the kernel functions particularly in the presence of missing data. A natural choice for $c_{1}$ and $c_{2}$ is based on the expected rate of decay over which the predictor is allowed to extend. Under this prior specification, a Metropolis-Hasting step is used to sample from the posterior distribution. Alternative choice of priors for different parameterizations of the scale parameters is discussed further in Section 4. Full details of the Metropolis-within-Gibbs sampler implementation are available in Appendix D.
3 Simulation Study
We conducted a simulation study to understand the behaviour of the model under different conditions. First we generated the spatio-temporal process via:
$\log(X_{t}(\mathbf{s}))=\mu_{t}(\mathbf{s})+\nu_{t}(\mathbf{s})$
with
$\mu_{t}(\mathbf{s})=$ $\exp\left(-\left(\frac{(s_{1}-s_{1,1,t})^{2}}{2\cdot 50^{2}}+\frac{(s_{2}-s_{2,1,t})^{2}}{2\cdot 10^{2}}\right)\right)+1.5\exp\left(-\left(\frac{(s_{1}-s_{1,2,t})^{2}}{2\cdot 70^{2}}+\frac{(s_{2}-s_{2,2,t})^{2}}{2\cdot 15^{2}}\right)\right)$
$+0.5\exp\left(-\left(\frac{(s_{1}-s_{1,3,t})^{2}}{2\cdot 40^{2}}+\frac{(s_{2}-s_{2,3,t})^{2}}{2\cdot 8^{2}}\right)\right)$
and $\nu_{t}(\mathbf{s})$ is a zero-mean spatial Gaussian field assuming an exponential covariance model $C(h)=\sigma^{2}\exp\left(-\frac{h}{\phi}\right)$ with variance $\sigma^{2}=1$ and scale (range) parameter $\phi=0.1$. Here $(s_{1,k,t},s_{2,k,t})$ denotes the centre of the $k$th (k=1,2,3) peak at time $t$; the three peak locations are redrawn at each time point. We then generate the data using
[…]
where $\varepsilon\sim\mathcal{N}(0,0.25^{2})$ and $c_{1}$ and $c_{2}$ are normalizing constants for the spatial and temporal kernels so that each integrates to 1 over its support. In this model specification, only predictors within a 10 km radius and the most recent 5 time lags affect the response (r=10,q=5r=10,\;q=5). Illustrations of the simulated surface and predictor process are provided in Appendix B. As shown in Figure 9, the spatio-temporal weights rapidly decay ($\to 0$) for $r>10$ and $l>4$.
We generated 100 replicated data sets on N = 100 locations and T = 153 time points. We retained N = 25 locations for model fitting and the remaining out-of-sample to compare predictive performance. This choice reflects the sparseness of ground-level stations in practice. Using the fully observed predictor process, we fit (i) a model with the true lag values $(r,q)=(10,5)$, (ii) a truncated misspecified model with $\tilde{r}=5<r$ and $\tilde{q}=2<q$ (iii) an extended misspecified model with $\tilde{r}=15>r$ and $\tilde{q}=8>q$. The covariate process $X_{t}(\mathbf{s})$ is well-behaved (no heavy tails or extremes). To assess the approximation under irregular sampling, we also simulated datasets with missing values that mimic the application’s missingness pattern observed (described below).For inference, we found that using a non-informative inverse-Gamma prior, $\sigma^{2}\sim\mathcal{IG}(a_{1},b_{1}),$ does not pose issues for parameter estimation or identifiability when the data are complete. However, in the presence of missing data, employing an informative prior helps improve both identifiability and sampler mixing.
Figure 1: Distribution of estimated model parameters across 100 simulated data sets with complete data (full) and missing data. Red boxplots indicate the results obtained using true radius and time lag values $(r=10,q=5)$, while blue $(\tilde{r}=15,\tilde{q}=8)$ and green $(\tilde{r}=5,\tilde{q}=2)$ illustrate the variability of estimates under misspecified radius and time lags.
Figure  1 illustrates the effect of model misspecification on parameter estimation. When $\tilde{r}<r$ and $\tilde{q}<q$, the model tends to overestimate parameters and exhibit greater variability than the correctly specified model. In particular, we observe large biases in the estimated scale parameter of the spatial kernel function. This pattern appeared both with full data setting and under missingness. A larger value or $\xi_{1}$ corresponds to a more uniform weighting across the radius (i.e. slower decay). Consequently, selecting a smaller radius can lead the model to overestimate the influence of predictors within $\tilde{r}$ to compensate for omitted contributions from $\mathcal{A}_{\tilde{r},r}(\mathbf{s})$. Similarly, the effect of parameter $\beta_{1}$ is overestimated to offset the mean bias induced by the reduced integration domain.
With missing data, the numerical approximation introduces some bias; even with the true $(r,q)$, parameter estimates are positively biased and the bias is typically exacerbated under model misspecification. The top left panel in Figure 2 displays the distribution of the noise parameter $\sigma^{2}$. Misspecification of the mean process results in systematically higher estimates of $\sigma^{2}$. In the presence of missing data, estimated $\sigma^{2}$ is generally larger, reflecting increased uncertainty.
Figure 2: Estimation of $\sigma^{2}$ and out-of-sample performance of simulated data measures using the $R^{2}$ metric, mean-squared error and coverage of the 95% prediction interval. Red boxplots indicate the results obtained using true radius and time lag values $(r=10,q=5)$, while blue $(\tilde{r}=15,\tilde{q}=8)$ and green $(\tilde{r}=5,\tilde{q}=2)$ illustrate the effect of model misspecification on predicitive performance.
Despite these biases, the $R^{2}$ values remain relatively high, particularly when a reduced region of integration is selected. This pattern persists in the presence of missing data. Similarly, the mean-squared error is low relative to the performance under the true parameters, although misspecified models exhibit greater variability in out-of-sample performance. Under the full data setting, coverage of the reduced model is lower, with prediction interval widths ranging from 2.81 to 6.30, compared to the wider intervals of the extended model (2.75 to 8.44). Finally, even with missing data, out-of-sample performance remains strong, with coverage exceeding 95%, reflecting wider prediction intervals and higher uncertainty estimates.
4 Application to PM2.5 mapping in Northern California
4.1 Data
The dataset comprises fine particulate matter (PM2.5) measurements from 67 air quality monitoring stations across Northern California between 1st June 2020 and 31st October 2020, coinciding with the region’s peak wildfire season. Ground-based data were augmented with satellite data of aerosol optical depth (AOD) from the Multi-Angle Implementation of Atmospheric Correction (MAIAC) algorithm, which provides high-resolution (1̃ km x 1 km) gridded observations from the Moderate Resolution Imaging Spectroradiometer (MODIS) instruments aboard NASA’s Terra and Aqua satellites [Lyapustin et al., 2018]. To account for wildfire smoke influence on PM2.5 concentrations, smoke plume data were obtained from the National Oceanic and Atmospheric Administration’s (NOAA) Hazard Mapping System (HMS), which delineates smoke plume extent based on satellite imagery [NOAA Satellite and Information Service, 2025]. Together these data enabled a comprehensive characterization of spatiotemporal variation in PM2.5 and its association with wildfire smoke during the 2020 fire season in Northern California.
Figure 3: Map of Northern California study region on two consecutive days in June 2020 showing: (a) PM2.5 monitoring locations (blue dots) with 20 km buffers around them (shaded blue circles) with 1 km x 1 km MAIAC aerosol optical depth (AOD) grid (pink to red); (b) zoom in of one PM2.5 monitor on a day with nearly completely observed AOD within 20 km; and (c) zoom in of one PM2.5 monitor on the next day with significant missing AOD within 20 km
4.2 Model Specifications and Estimation
In Table 1, we present the different models fitted to the data. The spatiotemporal baseline model assumes exponential kernels for both spatial and temporal lags. For the spatial kernel scale parameter we used a Gamma prior of the form $\mathcal{G}(a,b)$ where $a=1,b=9$ with the shape-scale parametrization $E(\xi_{1})=9$ placing mass on spatial scales consistent with AOD influence extending to 20 km. For the temporal kernel scale we used a weakly informative inverse-Gamma prior $\xi_{2}\sim\text{IG}(a_{0},b_{0})$.
Building on the baseline model, the functional mean model augments the intercept by a temporally varying mean function $\beta_{0}(t)$ represented with B-spline basis functions $f_{k}(t)$ [Wood, 2017]. We used $K=10$ equally spaced knots to capture bi-weekly trends, and retained the same priors as in the baseline model.
In the remaining models we included a varying coefficient for wildfire smoke. Let $I(\mathbf{s},t)$ indicate a smoke plume event (1 = satellite-derived wildfire smoke plume intersected a PM2.5 monitor $Y(\mathbf{s},t)$ at location $s$ on day $t$, 0 = otherwise). We incorporated this indicator to the models: (1) at the mean level where $\beta_{0}(\mathbf{s},t)=\sum_{k}b_{jk}f_{k}(t)+I(\mathbf{s},t)$, (2) as a varying coefficient where the effect of the weighted predictor of AOD varied by plume status $\beta_{1}(\mathbf{s},t)=\beta_{10}+\beta_{11}I(\mathbf{s},t)$, and (3) by fixing $\xi_{2}$ at the posterior median from the functional mean model and allowing the spatial scale to depend on smoke plume status. This was motivated by [Zhang, 2004], who showed that range and scale parameters in covariance functions are not consistently estimable and require stabilization.
The full model includes all features above and allows the spatial extent over which AOD influences PM2.5 to vary based on smoke plumes via $\log\{\xi_{1}(\mathbf{s},t)\}=\alpha_{0}+\alpha_{1}\,I(\mathbf{s},t)+\nu,\qquad\nu\sim N(0,\sigma_{\nu}^{2})$. To help estimation, when $I(\mathbf{s},t)=0$, we placed a strong point-mass prior on $\xi_{1}(\mathbf{s},t)$ centered at 2.0 for $r=10$ and 1.0 for $r=20$ (values obtained under the varying coefficient model). Under this approach, we effectively only estimated the effect of $I(\mathbf{s},t)$ through $\alpha_{1}$. The MCMC algorithms used for inference can be found under Appendix D. Convergence was assessed using the Gelman-Rubin test and effective sample size. Each model was run for 25000 MCMC iterations, with the first 12500 samples used as warm-up and retaining the latter 12500 of the samples for inference. As a benchmark, we also fit a collocated linear model where we aggregated and averaged AOD data within a 5 km radius around each station on each day.
Table 1: Spatio-temporal models for PM2.5: mean specification, temporal and spatial regression coefficients, and scale parameters.
Model $\beta_{0}(\mathbf{s},t)$ $\beta_{1}(\mathbf{s},t)$ $\xi_{1}(\mathbf{s},t)$ $\xi_{2}(\mathbf{s},t)$
Baseline $\beta_{0}$ $\beta_{1}$ $\xi_{1}\sim\mathrm{Gamma}(1,9)$ $\xi_{2}\sim\mathrm{IG}(a_{0},b_{0})$
Functional mean $\sum_{k}b_{0k}\,f_{k}(t)$ $\beta_{1}$ $\xi_{1}\sim\mathrm{Gamma}(1,9)$ $\xi_{2}\sim\mathrm{IG}(a_{0},b_{0})$
Varying coefficient $\sum_{k}b_{0k}\,f_{k}(t)$ $\gamma\,I(\mathbf{s},t)$ $\beta_{1}(\mathbf{s},t)=\beta_{10}$ $\beta_{11}I(\mathbf{s},t)$ $\xi_{1}\sim\mathrm{Gamma}(1,9)$ $\xi_{2}$ fixed at posterior median
Full $\sum_{k}b_{0k}\,f_{k}(t)$ $\gamma\,I(\mathbf{s},t)$ $\beta_{1}(\mathbf{s},t)=\beta_{10}$ $\beta_{11}I(\mathbf{s},t)$ $\log\!\{\xi_{1}(\mathbf{s},t)\}=\alpha_{0}$ $\alpha_{1}I(\mathbf{s},t)+\nu,$ $\nu\sim N(0,\sigma_{\nu}^{2})$ $\xi_{2}$ fixed at posterior median
Collocated Linear $\beta_{0}$ $\beta_{1}$
 Following [Zhang, 2004], $\xi_{2}$ is fixed at the posterior median from the functional mean model to stabilize range estimation. In all models, $\beta_{1}(\mathbf{s},t)$ multiplies $W(\mathbf{s},t)$, the kernel-weighted AOD predictor defined in Eq.(1), which averages $X(\mathbf{u},t-l)$ over $\mathcal{B}_{r}(\mathbf{s})\times[0,q]$ using normalized exponential kernels in space and time. The collocated baseline model uses the AOD average $\overline{X}_{5\text{km}}(\mathbf{s},t)$ in 5 km buffer instead.
We assumed independent Gaussian errors as our primary objective is to characterize the spatio-temporal dependence of PM2.5 on AOD. We intentionally excluded additional meteorological covariates to isolate this relationship and better understand the spatio-temporal patterns driven by AOD and wildfire smoke presence.
4.3 Results
The performance of each method is assessed using the root-mean square error (RMSE) metric, percentage of variance explained by the model (R2) and coverage of the 95% prediction interval for stations that were kept out-of-sample.
The baseline collocated linear model produced an RMSE of 25.8, $R^{2}=0.184$ and coverage 0.941. To fit this model, 407 observations were removed due to unpaired PM2.5-AOD observations.
Out-of-sample predictive performance of the competing models is reported in Table 2. Overall, the smaller buffer radius generally improved predictive accuracy, with the varying coefficient model benefitting less from this adjustment. In particular, the full model performs noticeably worse when the radius is set to $r=20$, suggesting that the added complexity at this radius may induce overfitting and reduce generalizability. However, we observed that the error is largely driven by overinflated predictions at a single location; when that location is removed, the model performance improves to RMSE = 19.3, R2 = 0.455, and coverage = 0.942.
By contrast, when the radius was restricted to $r=10$, the full model achieved the best balance across evaluation metrics, outperforming the alternative specifications. These findings highlight the sensitivity of model performance to the choice of buffer radius and emphasize the importance of carefully selecting this parameter in practice.
Table 2: Spatio-temporal model predictive performance for $r\in\{10,20\}$ km and $q=3$.
r=10,q=3r=10,\;q=3 r=20,q=3r=20,\;q=3
Model RMSE $\boldsymbol{R^{2}}$ Coverage RMSE $\boldsymbol{R^{2}}$ Coverage
Baseline 32.5 0.230 0.956 35.1 0.227 0.957
Functional mean 19.6 0.464 0.940 21.0 0.410 0.942
Varying coefficient 21.0 0.362 0.944 19.9 0.400 0.944
Full 19.0 0.454 0.940 63.3 0.037 0.940
Collocated baseline model: RMSE = 25.8, $R^{2}=0.184$, Coverage = 0.941.
To gain further insight into the spatio-temporal dependencies between PM2.5 and AOD, Figure 4 illustrates the estimated surface $\hat{\omega}(h,l)=\hat{\beta}_{1}\,\kappa_{1}(h;\hat{\xi}_{1}(\mathbf{s},t))\,\kappa_{2}(l;\hat{\xi}_{2}(\mathbf{s},t)),$ for the two best-performing models: the functional mean ST-Exp model (Figure 4(a)) and the full ST-Exp model (Figures 4(b) and 4(c)), both with $r=10$. We used the posterior medians of the MCMC samples to construct the surface $\widehat{\omega}(h,l)$. Figure 4(b) shows the estimated effect of AOD on PM2.5 when no smoke plume was present ($I(\mathbf{s},t)=0$), whereas Figure 4(c) shows the effect under plume conditions ($I(\mathbf{s},t)=1$). The results indicate that incorporating wildfire smoke information alters the spatio-temporal weighting: in the functional mean model, the AOD effect was distributed more evenly across the buffer region, while in the full model, both for $I(\mathbf{s},t)=0$ and $I(\mathbf{s},t)=1$, the effect decayed more rapidly over space and time but attained larger magnitudes in the immediate neighbourhood of an observation. Taken together, these results suggest that fire events amplify the local influence of AOD on PM2.5, underscoring the importance of accounting for such events when constructing accurate exposure maps.
(a) Functional mean
(b) Full model in the absence of fire plumes
(c) Full model in the presence of fire plumes
Figure 4: Estimates of the surface explaining the spatio-temporal dependencies between PM2.5 and AOD at varying spatial and time lags ($r=10$).
Figure 5: Predictions of PM2.5 at out-of-sample locations over time. The top panel compares the collocated linear model with the functional mean model ($r=10$). The bottom panel compares the functional mean model with the full model ($r=10$). As model complexity increases, uncertainty quantification of the predictions improves with narrower prediction intervals capturing peaks.
Figure 5 compares out-of-sample temporal predictions and 95% posterior prediction intervals at held-out monitoring stations. The collocated linear benchmark model attained reasonable point predictions but produced very wide intervals, especially during sharp concentration peaks. By contrast, the functional mean model has substantially narrower interval widths while maintaining coverage, providing more precise uncertainty quantification for PM2.5 concentrations across both peak episodes and stable periods. Figure 5(b) further shows that incorporating wildfire information improves temporal calibration. The full model produced narrower prediction intervals that still captured the observed peaks, highlighting the importance for wildfire influences in PM2.5 estimation.
Figure 6 presents 5 km resolution gridded maps of estimated PM2.5 derived from 1 km resolution AOD for three days in September during peak wildfire season, using the functional mean model. The approach effectively captured both spatial gradients and day-to-day temporal patterns, producing smooth, coherent variation across space and time, particularly where the August Complex wildfire occurred.
Figure 6: Spatial mapping of PM2.5 using the functional mean model for three consecutive days in September 2020.
5 Discussion
Handling spatial misalignment traditionally relies on change-of-support or hierarchical modeling to reconcile point- and areal-level data [Gotway and Young, 2002]. In point-to-point and point-to-grid misalignment, common practice is to average data from the misaligned source over buffer regions around a target point to align (i.e. forced collocation). This approach can clearly induce misclassification for several reasons including choice of the buffer radius. Furthermore, misalignment may be present due to missing data. Our contribution is to embed the missing data problem inside a normalized spatio-temporal weighting operator whose kernel parameters can vary with covariates, coupled with a numerically stable quadrature that respects missingness and irregular supports. This yielded a unified route to estimation and uncertainty quantification without a separate imputation stage, while remaining computationally tractable.
The weighted kernel-averaged predictor regression model we developed required careful specification of the spatial and temporal lag windows $(r,q)$ for valid inference and accurate prediction. Our theoretical results quantify how misspecification of $(r,q)$ propagated to bias in the mean, which in turn lead to biased parameter estimates and poorer predictive performance. Our proposed quadrature scheme that uses Voronoi-based spatial weights and irregular time steps to estimate the spatio-temporal integral, enabled estimation under extensive missing data and spatio-temporal misalignment. Unlike common strategies that first gap-fill (interpolate) and then regress on the complete interpolated surface (e.g. [Li and Heap, 2014]), our approach builds the weighting and uncertainty handling directly into the likelihood via $W(\mathbf{s},t)$, allowing propagation of uncertainty from the available data to the predictions.
Missingness and and spatial misalignment are common issues in many environmental modeling applications. In the satellite AOD–PM2.5 literature, missing AOD due to clouds, snow, high aerosol loadings, or retrieval failures is widespread, and many studies use explicit gap filling via interpolation or ensembles before regression [Xiao et al., 2021, Sorek-Hamer et al., 2020, Li et al., 2020]. When we applied our approach to estimate PM2.5 from satellite AOD over Northern California in 2020, which had an intense wildfire season, the weighted predictor model captured both spatial and temporal variation and yielded narrower, better-calibrated prediction intervals than a collocated linear baseline, with more than a 50% increase in R2. Incorporating smoke plume information altered the estimated weighting surface and improved temporal calibration around concentration peaks. The proposed estimator handled missing and misaligned data effectively because it aggregated over whatever nearby $(\mathbf{u},l)$ were observed, rather than requiring a fully gridded covariate. When no informative neighbours existed within $(r,q)$, the observation is flagged rather than imputed, avoiding potentially biased reconstructions.
Some limitations of our work remain. First, we focused on the PM2.5-AOD relationship and intentionally excluded meteorology to isolate this signal and develop methodology that targets the spatial misalignment and missing data issue. However, in practice models often adjust for other predictors to account for residual confounding from meteorological variables such as boundary-layer height, temperature, humidity, and winds. Thus, a natural extension to our work is to include spatio-temporally varying meteorology either as additional weighted predictors (with their own kernels and possibly interactions) or as modulators of the AOD kernel parameters $\boldsymbol{\xi}(\mathbf{s},t)$ (e.g., humidity-dependent temporal decay or wind-aligned spatial kernels). Second, we used exponential kernels; alternative families (e.g., Matern-like or compactly supported kernels) and nonseparable forms could better match plume transport physics. Third, for stability we fixed the temporal range in some specifications; fully hierarchical estimation of all kernel scales with stronger priors or penalization is an important avenue. Fourth, we assumed conditionally independent Gaussian errors; other residual structures (e.g., nonseparable errors) may further improve calibration, however may unnecessarily increase model complexity at the cost of computational intensity. Finally, scaling our models to larger spatial domains could computationally benefit from sparse quadrature, multi-resolution bases, or approximate inference (e.g. via INLA) tailored to kernel-weighted predictors.
Overall, the weighted predictor framework provides a robust and flexible tool for spatio-temporal exposure modeling beyond strictly concurrent effects, with particular strength under missing and misaligned covariates. Future work will integrate multi-source covariates (meteorology, emissions, chemical transport model outputs) and explore adaptive, plume-aware kernels to further enhance predictive performance and interpretability.
Appendix A Voronoi Tessellation
Figure 7: Example of a Voronoi tessellation over a buffer.
Appendix B Simulation Study
Figure 8: Examples of simulated $X(s,t)$ fields.
Figure 9: Examples of weight surface used in simulation study.
Appendix C Proof of Proposition 1
Proof.
We consider the four possible cases.
  1. (1)
    $\tilde{r}\neq r$ and $\tilde{q}=q$. The two spatial balls differ, but the time window is the same. The symmetric–difference region is the annulus
[…]
    Thus
    $\bigl|\mu(\mathbf{s},t)-\tilde{\mu}(\mathbf{s},t)\bigr|\leq\beta\int_{0}^{q}\int_{\mathcal{A}_{r,\tilde{r}}(\mathbf{s})}\kappa_{1}(\mathbf{s}-\mathbf{u})\,\kappa_{2}(t-\ell)\,X(\mathbf{u},t-\ell)\,d\mathbf{u}\,d\ell.$
  2. (2)
    $\tilde{q}\neq q$ and $\tilde{r}=r$. The spatial region is the same, but the time windows differ. The symmetric–difference region is the extra time slab
[…]
    over the common ball $\mathcal{B}_{r}(\mathbf{s})$. Hence
    $\bigl|\mu(\mathbf{s},t)-\tilde{\mu}(\mathbf{s},t)\bigr|\leq\beta\int_{\mathcal{B}_{r}(\mathbf{s})}\int_{\min(q,\tilde{q})}^{\max(q,\tilde{q})}\kappa_{1}(\mathbf{s}-\mathbf{u})\,\kappa_{2}(t-\ell)\,X(\mathbf{u},t-\ell)\,d\ell\,d\mathbf{u}.$
  3. (3)
    $\tilde{r}>r$ and $\tilde{q}>q$. The region included by the larger window consists of
[…]
    The bounding region is
[…]
    Since $A\subset B$ and $\kappa_{1},\kappa_{2},X$ are nonnegative, the integral over $A$ is bounded by the integral over $B$, giving the desired inequality.
  4. (4)
    $\tilde{r}<r$ and $\tilde{q}<q$. The argument is analogous. The omitted region is
    $A=\bigl(\mathcal{A}_{\tilde{r},r}(\mathbf{s})\times[0,\tilde{q}]\bigr)\;\cup\;\bigl(\mathcal{B}_{r}(\mathbf{s})\times[\tilde{q},q]\bigr),$
    and the bounding region
    $B=\bigl(\mathcal{A}_{\tilde{r},r}(\mathbf{s})\times[0,q]\bigr)\;\cup\;\bigl(\mathcal{B}_{r}(\mathbf{s})\times[\tilde{q},q]\bigr)$
    satisfies $A\subset B$. Nonnegativity of the integrand again yields the bound.
  5. (5)
    $\tilde{r}>r$, $\tilde{q}<q$. The mismatch arises from a larger spatial radius but a smaller temporal lag.
    (i) The part omitted by the shorter temporal window is
    $A_{1}=\mathcal{B}_{r}(\mathbf{s})\times[\tilde{q},\,q].$
    (ii) The part added by enlarging the spatial radius is the annular region
    $A_{2}=\mathcal{A}_{r,\tilde{r}}(\mathbf{s})\times[0,\,\tilde{q}$].
    Thus the absolute difference of the two integrals equals the integral of the (nonnegative) integrand over $A_{2}$ minus the integral over $A_{1}$ in magnitude, and is therefore bounded above by the integral of the integrand over
    $A_{1}\cup A_{2}.$
    The given bounding region is
    $B=\bigl(\mathcal{A}_{r,\tilde{r}}(\mathbf{s})\times[0,\,q]\bigr)\;\cup\;\bigl(\mathcal{B}_{\tilde{r}}(\mathbf{s})\times[\tilde{q},\,q]\bigr).$
    Since $A_{1}\cup A_{2}\subset B$ and the integrand $\kappa_{1}\,\kappa_{2}\,X$ is nonnegative, the difference $|\mu(\mathbf{s},t)-\tilde{\mu}(\mathbf{s},t)|$ is bounded by the integral over $B$.
  6. (6)
    $\tilde{r}<r$, $\tilde{q}>q$. The mismatch arises from a larger temporal lag but a smaller spatial radius.
    (i) The part omitted by the smaller radius is
    $A_{1}=\mathcal{A}_{\tilde{r},r}(\mathbf{s})\times[0,\,q].$
    (ii) The part added by increasing temporal lag
    $A_{2}=\mathcal{B}_{\tilde{r}}(\mathbf{s})\times[q,\tilde{q}$].
    Thus the absolute difference of the two integrals equals the integral of the (nonnegative) integrand over $A_{1}$ minus the integral over $A_{2}$ in magnitude, and is therefore bounded above by the integral of the integrand over
    $A_{1}\cup A_{2}.$
    The given bounding region is
[…]
    Since $A_{1}\cup A_{2}\subset B$ and the integrand $\kappa_{1}\,\kappa_{2}\,X$ is nonnegative, the difference $|\mu(\mathbf{s},t)-\tilde{\mu}(\mathbf{s},t)|$ is bounded by the integral over $B$.
Appendix D Brief description of MCMC algorithm
Under a Gaussian likelihood, the conditional posterior distribution $\beta|\cdot$ is given by
$\mathbf{\beta}|\mathbf{\xi},\sigma^{2}\propto\pi(\mathbf{\beta})\times\prod_{i,j}f(Y(s_{i},t_{j})\;;W(s_{i},t_{j}\;;\xi(s_{i},t_{j})),\sigma^{2})$ (A1.1)
where $f(\cdot\;;\mu,\sigma^{2})$ is the p.d.f of a Gaussian distribution. Under a Gaussian prior distribution $\mathcal{N}(0,\sigma_{b}^{2})$, $\mathbf{\beta}|\cdot$ is conditionally normally distributed. We assume a conditionally conjugate inverse-gamma prior for $\sigma_{b}^{2}$,
$\sigma_{b}^{2}|\cdot\propto\pi(\sigma_{b})\times f(\beta\;;0,\sigma_{b}^{2})$ (A1.2)
Similarly, the conditional posterior distribution $\sigma^{2}|\cdot$ is given by,
$\mathbf{\sigma}^{2}|\mathbf{\xi},\mathbf{\beta}\propto\pi(\sigma^{2})\times\prod_{i,j}f(Y(s_{i},t_{j})\;;W(s_{i},t_{j}\;;\xi(s_{i},t_{j})),\sigma^{2}).$ (A2)
We assign the prior distribution $\sigma^{2}\sim\mathcal{IG}(0.001,0.001)$ which results in a conditional posterior distribution for $\sigma^{2}|\cdot$ that also follows an inverse-gamma distribution.
1. For the baseline, functional mean and varying coefficient models where $\xi_{1}(\mathbf{s},t)=\xi_{1}$ and $\xi_{2}(\mathbf{s},t)=\xi_{2}$, we assume a Gamma prior distribution, which leads to the conditional posterior distribution for $\xi_{1}|\cdot$ and $\xi_{2}|\cdot$,
$\mathbf{\xi}_{1}|\cdot$ $\propto\pi(\xi_{1})\times\prod_{i,j}f(Y(s_{i},t_{j})\;;W(s_{i},t_{j}\;;\xi(s_{i},t_{j})),\sigma^{2})$ (A3)
$\mathbf{\xi}_{2}|\cdot$ $\propto\pi(\xi_{2})\times\prod_{i,j}f(Y(s_{i},t_{j})\;;W(s_{i},t_{j}\;;\xi(s_{i},t_{j})),\sigma^{2})$
which are intractable.
2. For the full model where $\log\{\xi_{1}(\mathbf{s},t)\}=\alpha_{0}+\alpha_{1}+I(\mathbf{s},t)+\nu$ we assume (i) a point-mass prior distribution when $I(\mathbf{s},t)=0$, $\log\{\xi_{1}(\mathbf{s},t)\}=\hat{\xi}_{1,I=0}$, (ii) when $I(\mathbf{s},t)=1$, $\log\{\xi_{1}(\mathbf{s},t)\}=\log(\xi_{1,I=1})=\alpha_{0}+\alpha_{1}+\nu,\;\nu\sim\mathcal{N}(0,\sigma^{2}_{\nu}).$
  1. 1.
    The conditional posterior distribution for $\xi_{1,I=1}$ is given by
    $\mathbf{\xi}_{1,I=1}|\cdot$ $\propto f(\log(\xi_{1,I=1})\;;\alpha_{0}+\alpha_{1},\sigma^{2}_{\alpha})\times\prod_{i,j}f(Y(s_{i},t_{j})\;;W(s_{i},t_{j}\;;\xi(s_{i},t_{j}))$ (A4)
  2. 2.
    Assuming a bivariate Gaussian prior distribution for $\mathbf{\alpha}$,
    $\mathbf{\alpha}|\cdot\propto f(\mathbf{\alpha}\;;\mathbf{\mu}_{\alpha},\sigma^{2}_{\alpha}\mathbf{I})\times\prod_{i,t}f(\log(\xi(s_{i},t_{j}))\;;\alpha_{0}+\alpha_{1},\sigma_{\nu}^{2})$ (A5)
    which leads to a conditionally Gaussian posterior distribution.
  3. 3.
    Assuming an inverse-gamma prior distribution for $\sigma_{\nu}^{2}$,
    $\sigma_{\nu}^{2}|\cdot\propto\mathcal{\pi}(\sigma_{\nu}^{2})\times\prod_{i,t}f(\log(\xi(s_{i},t_{j}))\;;\alpha_{0}+\alpha_{1},\sigma_{\nu}^{2})$ (A6)
    which leads to an inverse-gamma conditional posterior distribution.
The MCMC algorithms are given in the tables below:
for $t\leftarrow 1$ to $T$ do
 1. Sample $\mathbf{\beta}^{(t)}$ from $\mathbf{\beta}\mid\cdot$ (A1.1).;
  2mm 2. Sample $\sigma_{b}^{2^{(t)}}$ from $\sigma_{b}^{2}\mid\cdot$ (A1.2).;
  2mm 3. For each $m\in\{1,2\}$;
   (random-walk Metropolis on log-scale):;
  1. 1.
    Draw proposal on the log-scale
    $\log(\xi_{m}^{\prime})\sim\mathcal{N}\big(\log(\xi_{m}^{(t-1)}),\,\delta_{1}\big),$
    and set $\xi_{m}^{\prime}=\exp(\log(\xi_{m}^{\prime}))$.
  2. 2.
    Compute the Metropolis–Hastings acceptance probability using (A4):
    $p\;=\;\min\!\left\{1,\;\frac{\pi(\xi_{m}^{\prime}\mid\text{rest})}{\pi(\xi_{m}^{(t-1)}\mid\text{rest})}\cdot\frac{q(\xi_{m}^{(t-1)}\mid\xi_{m}^{\prime})}{q(\xi_{m}^{\prime}\mid\xi_{m}^{(t-1)})}\right\}.$
    Because the proposal is Gaussian on the log scale,
    $\frac{q(\xi_{m}^{(t-1)}\mid\xi_{m}^{\prime})}{q(\xi_{m}^{\prime}\mid\xi_{m}^{(t-1)})}\;=\;\frac{1/\xi_{m}^{(t-1)}}{1/\xi_{m}^{\prime}}\;=\;\frac{\xi_{m}^{\prime}}{\xi_{m}^{(t-1)}}.$
    Hence the working form is
    $p\;=\;\min\!\left\{1,\;\frac{\pi(\xi_{m}^{\prime}\mid\text{rest})}{\pi(\xi_{m}^{(t-1)}\mid\text{rest})}\cdot\frac{\xi_{m}^{\prime}}{\xi_{m}^{(t-1)}}\right\}.$
  3. 3.
    Draw $u\sim\mathcal{U}(0,1)$. If $u\leq p$ then accept: $\xi_{m}^{(t)}\leftarrow\xi_{m}^{\prime}$, else reject: $\xi_{m}^{(t)}\leftarrow\xi_{m}^{(t-1)}$, and update $W(\mathbf{s},\tau;\mathbf{\xi}^{(t)})\leftarrow W(\mathbf{s},\tau;\mathbf{\xi}^{(t-1)})$.
  2mm
 4. Sample $\sigma^{{}^{2}{(t)}}$ from $\sigma^{2}\mid\cdot$ (A2)
end for
Algorithm 1 Metropolis-within-Gibbs when $\xi_{m}(\mathbf{s},t)=\xi_{m}$
for $t\leftarrow 1$ to $T$ do
 1. Sample $\mathbf{\beta}^{(t)}$ from $\mathbf{\beta}\mid\cdot$ (A1.1).;
  2mm 2. Sample $\sigma_{b}^{2^{(t)}}$ from $\sigma_{b}^{2}\mid\cdot$ (A1.2).;
  2mm
 3. Fix $\xi_{2}$ and $\xi_{1}(s,t)$ when $I(s,t)=0$. For $\xi_{1}(s,t)=\xi_{1,I=1}$ when $I(s,t)=1$ ;
   (random-walk Metropolis on log-scale):;
  1. 1.
    Draw proposal on the log-scale
    $\log(\xi_{1,I=1}^{\prime})\sim\mathcal{N}\big(\log(\xi_{1,I=1}^{(t-1)}),\,\delta_{1}\big),$
    and set $\xi_{1,I=1}^{\prime}=\exp(\log(\xi_{1,I=1}^{\prime}))$.
  2. 2.
    Compute the Metropolis–Hastings acceptance probability using (A4):
    $p\;=\;\min\!\left\{1,\;\frac{\pi(\xi_{1,I=1}^{\prime}\mid\text{rest})}{\pi(\xi_{1,I=1}^{(t-1)}\mid\text{rest})}\cdot\frac{q(\xi_{1,I=1}^{(t-1)}\mid\xi_{m}^{\prime})}{q(\xi_{1,I=1}^{\prime}\mid\xi_{1,I=1}^{(t-1)})}\right\}.$
    Because the proposal is Gaussian on the log scale,
    $\frac{q(\xi_{1,I=1}^{(t-1)}\mid\xi_{1,I=1}^{\prime})}{q(\xi_{1,I=1}^{\prime}\mid\xi_{1,I=1}^{(t-1)})}\;=\;\frac{1/\xi_{1,I=1}^{(t-1)}}{1/\xi_{1,I=1}^{\prime}}\;=\;\frac{\xi_{1,I=1}^{\prime}}{\xi_{1,I=1}^{(t-1)}}.$
    Hence the working form is
    $p\;=\;\min\!\left\{1,\;\frac{\pi(\xi_{1,I=1}^{\prime}\mid\text{rest})}{\pi(\xi_{1,I=1}^{(t-1)}\mid\text{rest})}\cdot\frac{\xi_{1,I=1}^{\prime}}{\xi_{1,I=1}^{(t-1)}}\right\}.$
  3. 3.
    Draw $u\sim\mathcal{U}(0,1)$. If $u\leq p$ then accept: $\xi_{1,I=1}^{(t)}\leftarrow\xi_{1,I=1}^{\prime}$, else reject: $\xi_{1,I=1}^{(t)}\leftarrow\xi_{1,I=1}^{(t-1)}$, and update $W(\mathbf{s},\tau;\mathbf{\xi}^{(t)})\leftarrow W(\mathbf{s},\tau;\mathbf{\xi}^{(t-1)})$.
  2mm
 4. Sample $\sigma^{2^{(t)}}$ from $\sigma^{2}\mid\cdot$ (A2)
 5. Sample $\mathbf{\alpha}^{{(t)}}$ from $\mathbf{\alpha}\mid\cdot$ (A5)
 6. Sample $\mathbf{\sigma}_{\nu}^{2^{(t)}}$ from $\sigma_{\nu}^{2}\mid\cdot$ (A6)
end for
Algorithm 2 Metropolis-within-Gibbs when $\log(\xi_{m}(\mathbf{s},t))=\alpha_{0}+\ \alpha_{1}I(s,t)+\nu$

### v2

@@
-Bayesian spatio-temporal weighted regression for integrating missing and misaligned environmental data
+Methodology, Formal Analysis, Software, Validation, Visualization, Writing – original draft
 Abstract
@@
 We showed that misspecification of the spatial and temporal lags induced bias in the mean and parameter estimates, indicating the need for principled parameter selection.
-Simulation studies confirmed these theoretical findings, where careful tuning was critical to control bias and achieve accurate prediction, while the proposed quadrature performed well under severe missingness.
+Simulation studies confirmed these findings, where careful tuning was critical to control bias and achieve accurate prediction, while the proposed quadrature performed well under severe missingness.
 As an illustrative application, we estimated fine particulate matter (PM2.5) in northern California using satellite-derived aerosol optical depth (AOD) and wildfire smoke plume indicators.
-Relative to a traditional collocated linear model, our approach improved out-of-sample predictive performance (over 50% increase in R2), reduced uncertainty, and yielded robust temporal predictions and spatial surface estimation.
+Relative to a traditional collocated linear model, our approach improved out-of-sample predictive performance, reduced uncertainty, and yielded robust temporal predictions and spatial surface estimation.
 Our framework is extensible to additional spatio-temporally varying covariates and other kernel families.
-Keywords Spatio-temporal regression $\cdot$ Kernel-averaged predictor $\cdot$ Missing and misaligned data $\cdot$ PM2.5 mapping
+keywords:
+[…]
 1 Introduction
 Environmental exposures derived from multi-source data play a central role in public health.
-Accurate exposure estimates are needed to quantify population-level risks, guide interventions, and inform regulatory action [Clark et al., 2025].
-Data streams such as satellite observations, ground monitors, and physical model outputs provide complementary information, but they often differ in spatial resolution, temporal frequency, and coverage [Lolli, 2025].
-Integrating these sources is essential for many applications but can be methodologically challenging [Forlani et al., 2020].
+Accurate exposure estimates are needed to quantify population-level risks, guide interventions, and inform regulatory action (clark2025).
+Data streams such as satellite observations, ground monitors, and physical model outputs provide complementary information, but they often differ in spatial resolution, temporal frequency, and coverage (lolli2025).
+Integrating these sources is essential for many applications but can be methodologically challenging (forlani2020).
 Missing observations, spatial or temporal misalignment, and nonstationary associations between predictors and outcomes introduce uncertainty, limiting predictive reliability and the robustness of scientific inference.
 Air quality research provides a clear example of these challenges.
-Exposure to fine particulate matter (PM2.5) is a leading risk factor for respiratory and cardiovascular disease and premature mortality worldwide, contributing to an estimated 8.1 million deaths globally in 2021 [Health Effects Institute, 2024].
-While ground-based monitoring networks provide reliable measurements, they are spatially sparse and unevenly distributed [Zhang et al., 2021], limiting their ability to capture fine-scale spatial variability.
-Satellite-derived aerosol optical depth (AOD), available from instruments like the Moderate resolution Imaging Spectroradiometer (MODIS), offer global spatial coverage and have been widely used as proxies for surface PM2.5 [Diao et al., 2019].
+Exposure to fine particulate matter (PM2.5) is a leading risk factor for respiratory and cardiovascular disease and premature mortality worldwide, contributing to an estimated 8.1 million deaths globally in 2021 (health_effects_institute_state_2024).
+While ground-based monitoring networks provide reliable measurements, they are spatially sparse and unevenly distributed (zhang_satellite_2021), limiting their ability to capture fine-scale spatial variability.
+Satellite-derived aerosol optical depth (AOD), available from instruments like the Moderate resolution Imaging Spectroradiometer (MODIS), offer global spatial coverage and have been widely used as proxies for surface PM2.5 (diao2019).
 However, AOD retrievals are often missing due to cloud cover, snow (bright surfaces), or algorithmic failures, and when available, they are not always co-located or spatially aligned with ground monitors, leading to incomplete or misaligned data.
 To address these limitations, statistical approaches have been developed to model the PM2.5–AOD relationship and integrate additional sources of data such as meteorology and land use characteristics.
-Early methods relied on linear regression [Wang and Christopher, 2003, Chu et al., 2003] and were later extended through multiple linear regression and generalized additive models to incorporate meteorological covariates [van Donkelaar et al., 2006, Liu et al., 2009, Franklin et al., 2017].
+Early methods relied on linear regression (wang_intercomparison_2003; chu_global_2003) and were later extended through multiple linear regression and generalized additive models to incorporate meteorological covariates (van_donkelaar_estimating_2006; liu_estimating_2009; franklin2017).
 In these studies, spatial misalignment was commonly addressed by interpolating or averaging AOD onto regular grids first, then modeling the PM-AOD relationship.
-Machine learning and deep learning have come to the forefront in interpolation efforts as well, but explicit spatio-temporal trends are not generally incorporated in these approaches [Li et al., 2020].
-Multi-stage spatio-temporal models have been proposed to estimate daily PM2.5 using neighbouring observations where AOD is missing, [Pu and Yoo, 2020]; however, these approaches remain limited in their ability to capture nonstationarity and locally varying relationships.
+Machine learning and deep learning have come to the forefront in interpolation efforts as well, but explicit spatio-temporal trends are not generally incorporated in these approaches (li2020).
+Multi-stage spatio-temporal models have been proposed to estimate daily PM2.5 using neighbouring observations where AOD is missing, (pu_spatio-temporal_2020); however, these approaches remain limited in their ability to capture nonstationarity and locally varying relationships.
 Local regression techniques provide one pathway to address this gap.
-Geographically weighted regression (GWR) [Brunsdon et al., 1998] and its spatio-temporal extension, GTWR [Fotheringham et al., 2015] apply location- and time-specific weights, allowing regression coefficients to vary across both space and time.
-GTWR has been shown to improve predictive performance for PM2.5 mapping by capturing nonstationary AOD–PM2.5 relationships [He and Huang, 2018, Hu et al., 2013], producing fine-resolution exposure surfaces that reveal gradients and hotspots.
+Geographically weighted regression (GWR) (brunsdon_geographically_1998) and its spatio-temporal extension, GTWR (fotheringham_geographical_2015) apply location- and time-specific weights, allowing regression coefficients to vary across both space and time.
+GTWR has been shown to improve predictive performance for PM2.5 mapping by capturing nonstationary AOD–PM2.5 relationships (he_satellite-based_2018; hu_estimating_2013), producing fine-resolution exposure surfaces that reveal gradients and hotspots.
 Yet, these models still rely on concurrent predictor–outcome pairs, making them vulnerable to the systematic missingness and misalignment common in environmental datasets.
 Other methodological advances in spatial modeling have aimed to extend beyond strictly concurrent data.
-Kernel-averaged predictor frameworks [Heaton and Gelfand, 2011, Heaton and Gelfand, 2012] smoothed predictors over a spatial or spatio-temporal buffer, allowing nearby information to inform estimates even when direct matches are unavailable.
+Kernel-averaged predictor frameworks (heaton_spatial_2011; heaton_kernel_2012) smoothed predictors over a spatial or spatio-temporal buffer, allowing nearby information to inform estimates even when direct matches are unavailable.
 While promising, these methods can be sensitive to the choice of kernel size and lag parameters, and their robustness to missing data remains unclear.
 This is a pressing concern in many applications, where monitoring networks are sparse, satellite retrievals are often intermittent, and episodic events such as wildfire smoke plumes create rapidly evolving exposure patterns.
-In this paper, we build on the developments of [Heaton and Gelfand, 2012] by proposing a new method of estimation for the weighted predictor regression framework, tailored to settings with missing or misaligned data.
+In this paper, we build on the developments of heaton_kernel_2012 by proposing a new method of estimation for the weighted predictor regression framework, tailored to settings with missing or misaligned data.
 Our Bayesian formulation allows flexible spatial and temporal lag structures, enabling regression even when predictors are not observed concurrently with the outcome.
@@
 We denote a spatio-temporal process by $\{Y(\mathbf{s},t):\mathbf{s}\in\mathbb{R}^{d},t\in\mathbb{R}\}$ where $Y(\mathbf{s},t)$ represents an observation at location $\mathbf{s}$ and time $t$ and assume that $Y(\mathbf{s},t)$ depends on another spatio-temporal temporal process $\{X(\mathbf{u},\tau):\mathbf{u}\in\mathbb{R}^{d},\tau\in\mathbb{R}\}$.
-In a general setting, the proposed model, similar to [Heaton and Gelfand, 2012], is given by
+In a general setting, the proposed model, similar to heaton_kernel_2012, is given by
 $\mathrm{Y}(\mathbf{s},t)=\beta_{0}(\mathbf{s},t)+\beta_{1}(\mathbf{s},t)\mathrm{W}(\mathbf{s},t)+\varepsilon(\mathbf{s},t).$ (1)
@@
 This framework captures spatio-temporal variation by accounting for the local weighting of nearby predictors and recent time points, as well as the regression effect.
-These two sets of parameters play different roles: the (kernel) decay parameters $\xi(\mathbf{s},t)$ determine the spatial extent over which the predictors can influence the parameters, whereas the regression parameters $\beta_{j}(\mathbf{s},t)$ govern the strength of the association between the weighted predictor and the response.
+These two sets of parameters play different roles: the (kernel) decay parameters $\xi(\mathbf{s},t)$ determine the spatial extent over which the predictors can influence the parameters, whereas the regression parameters $\beta_{1}(\mathbf{s},t)$ govern the strength of the association between the weighted predictor and the response.
 We model both sources of non-stationarity using smooth functions or covariates observed at the corresponding locations and times:
-$\beta_{j}(\mathbf{s},t)=\sum b_{jk}f_{k}(s,t)\textrm{ and }\mathbb{E}(\xi(\mathbf{s},t))=\sum_{g=1}^{G}e_{g}v_{g}(s,t),\;\xi(\mathbf{s},t)\sim\mathcal{G}(\cdot)$
+$\beta_{1}(\mathbf{s},t)=\sum b_{k}f_{k}(s,t)\textrm{ and }\mathbb{E}(\xi(\mathbf{s},t))=\sum_{g=1}^{G}e_{g}v_{g}(s,t),\;\xi(\mathbf{s},t)\sim\mathcal{F}(\cdot)$
 where $f_{k}$ and $v_{g}$ represent smooth basis functions (e.g.
@@
 The choice of $r$ and $q$ must be treated as a model selection problem in which values that maximize goodness-of-fit measures such as R2 and/or minimize Akaike Information Criteria (AIC) and mean-squared error (MSE) are selected.
-[Heaton and Gelfand, 2011] choose $r$ based on the effective spatial range of the predictor.
+heaton_spatial_2011 choose $r$ based on the effective spatial range of the predictor.
 However, the effect of misspecifying $r$ and $q$ on parameter estimation and predictive performance remains unexplored.
@@
 Proposition 1 implies that the discrepancy between the true and misspecified means is influenced not only by the behaviour of the kernel functions but also by the magnitude of the process outside the true spatial–temporal window.
+The proof of this proposition can be found under Appendix C of the Supplementary Material.
 Corollary 1.
@@
 For example, on a regular 1 km² daily grid, the number of data points required to cover a region of radius $\mathcal{B}(\tilde{r})\times[0,\tilde{q}$] is $\pi\tilde{r}^{2}\tilde{q}$, which grows quadratically with $\tilde{r}$.
-Thus, it is recommended that model selection be based on performance assessed under a set of varying spatial and temporal lag values.
+Thus, it is recommended that model selection be based on out-of-sample performance assessed under a set of varying spatial and temporal lag values.
 2.2 Model estimation
@@
 When data are missing or misaligned, this estimation is more challenging.
-[Heaton and Gelfand, 2011] use Monte Carlo integration and [Heaton and Gelfand, 2012] project $X(\mathbf{s},t)$ onto a regular grid and approximate the integral by summing pointwise products of the discretised covariate with its weights after fitting a latent Gaussian process - an approach that can be computationally expensive for finely spaced data due to covariance estimation.
+heaton_spatial_2011 use Monte Carlo integration and heaton_kernel_2012 project $X(\mathbf{s},t)$ onto a regular grid and approximate the integral by summing pointwise products of the discretised covariate with its weights after fitting a latent Gaussian process - an approach that can be computationally expensive for finely spaced data due to covariance estimation.
 We instead use a numerical approximation for the weighted predictor function:;
@@
 Suppose $X(s_{k},\cdot)$ is observed at time-points $(t_{1},\ldots,t_{k-1},t_{k},\ldots,T_{X_{k}})$ then $\delta_{k}=t_{k}-t_{k-1}$.
-A Voronoi tessellation is a computational technique that partitions a metric space based on proximity to a given set of points [Okabe et al., 2009].
+A Voronoi tessellation is a computational technique that partitions a metric space based on proximity to a given set of points okabe_spatial_2009.
 At time $i$, given a finite set of distinct points given by $\mathbf{z}_{ni}$ which represents the locations at which we have observed data within the buffer of location $n$, the Voronoi cell $V_{ni,k}$ associated with location $z_{ni,k}$ is defined as:
@@
 In this case, $Y(\mathbf{s}_{n},t_{i})$ should be removed to avoid large approximation error.
-In Appendix A Figure 7, we illustrate the Voronoi tessellation of the buffer on two different days.
+In Appendix A, we illustrate the Voronoi tessellation of the buffer on two different days.
 The tessellation depends on the location of the observed AOD data.
+An advantage of this method is that it accounts for regions where AOD is unobserved by approximating them using the nearest observed values, rather than treating them as missing, as is often assumed in traditional collocation approaches.
 2.2.2 Bayesian inference
 A Metropolis-within-Gibbs sampler is used to perform inference.
-We assume a Gaussian for prior $\mathbf{\beta}\sim\mathcal{N}(0,\sigma^{2}_{b})$.
+We assume a Gaussian for prior $b_{k}\sim\mathcal{N}(0,\sigma^{2}_{b})$.
 This choice of prior is conditionally conjugate under a Gaussian process assumption for $\varepsilon(\mathbf{s},t)\sim\mathcal{GP}(0,\sigma^{2}C(\mathbf{s},\mathbf{s}^{\prime},t,t^{\prime}))$.
@@
 with
-$\mu_{t}(\mathbf{s})=$ $\exp\left(-\left(\frac{(s_{1}-s_{1,1,t})^{2}}{2\cdot 50^{2}}+\frac{(s_{2}-s_{2,1,t})^{2}}{2\cdot 10^{2}}\right)\right)+1.5\exp\left(-\left(\frac{(s_{1}-s_{1,2,t})^{2}}{2\cdot 70^{2}}+\frac{(s_{2}-s_{2,2,t})^{2}}{2\cdot 15^{2}}\right)\right)$
-$+0.5\exp\left(-\left(\frac{(s_{1}-s_{1,3,t})^{2}}{2\cdot 40^{2}}+\frac{(s_{2}-s_{2,3,t})^{2}}{2\cdot 8^{2}}\right)\right)$
+$\mu_{t}(\mathbf{s})=$ $\exp\left(-\left(\frac{(s_{1}-s_{1,1,t})^{2}}{2\cdot 50^{2}}+\frac{(s_{2}-s_{2,1,t})^{2}}{2\cdot 10^{2}}\right)\right)+1.5\exp\left(-\left(\frac{(s_{1}-s_{1,2,t})^{2}}{2\cdot 70^{2}}+\frac{(s_{2}-s_{2,2,t})^{2}}{2\cdot 15^{2}}\right)\right)+$ (4)
+$0.5\exp\left(-\left(\frac{(s_{1}-s_{1,3,t})^{2}}{2\cdot 40^{2}}+\frac{(s_{2}-s_{2,3,t})^{2}}{2\cdot 8^{2}}\right)\right)$ (5)
 and $\nu_{t}(\mathbf{s})$ is a zero-mean spatial Gaussian field assuming an exponential covariance model $C(h)=\sigma^{2}\exp\left(-\frac{h}{\phi}\right)$ with variance $\sigma^{2}=1$ and scale (range) parameter $\phi=0.1$.
@@
 In this model specification, only predictors within a 10 km radius and the most recent 5 time lags affect the response (r=10,q=5r=10,\;q=5).
-Illustrations of the simulated surface and predictor process are provided in Appendix B.
-As shown in Figure 9, the spatio-temporal weights rapidly decay ($\to 0$) for $r>10$ and $l>4$.
+As shown in the illustrations of the simulated surface and predictor process provided in Appendix B, the spatio-temporal weights rapidly decay ($\to 0$) for $r>10$ and $q>3$.
+Finally, we set $\beta_{0}=10$ and $\beta_{1}=0.5$.
 We generated 100 replicated data sets on N = 100 locations and T = 153 time points.
@@
 Using the fully observed predictor process, we fit (i) a model with the true lag values $(r,q)=(10,5)$, (ii) a truncated misspecified model with $\tilde{r}=5<r$ and $\tilde{q}=2<q$ (iii) an extended misspecified model with $\tilde{r}=15>r$ and $\tilde{q}=8>q$.
-The covariate process $X_{t}(\mathbf{s})$ is well-behaved (no heavy tails or extremes).
-To assess the approximation under irregular sampling, we also simulated datasets with missing values that mimic the application’s missingness pattern observed (described below).For inference, we found that using a non-informative inverse-Gamma prior, $\sigma^{2}\sim\mathcal{IG}(a_{1},b_{1}),$ does not pose issues for parameter estimation or identifiability when the data are complete.
+The covariate process $X_{t}(\mathbf{s})$ is well-behaved (no extremes).
+To assess the approximation under irregular sampling, we also simulated datasets with missing values that mimic the application’s missingness pattern observed (described in Figure 3).
+For inference, we found that using a non-informative inverse-Gamma prior, $\sigma^{2}\sim\mathcal{IG}(a_{1},b_{1}),$ does not pose issues for parameter estimation or identifiability when the data are complete.
 However, in the presence of missing data, employing an informative prior helps improve both identifiability and sampler mixing.
@@
 Figure  1 illustrates the effect of model misspecification on parameter estimation.
-When $\tilde{r}<r$ and $\tilde{q}<q$, the model tends to overestimate parameters and exhibit greater variability than the correctly specified model.
+When $\tilde{r}<r$ and $\tilde{q}<q$, the model is more prone to overestimating parameters.
 In particular, we observe large biases in the estimated scale parameter of the spatial kernel function.
@@
 Consequently, selecting a smaller radius can lead the model to overestimate the influence of predictors within $\tilde{r}$ to compensate for omitted contributions from $\mathcal{A}_{\tilde{r},r}(\mathbf{s})$.
-Similarly, the effect of parameter $\beta_{1}$ is overestimated to offset the mean bias induced by the reduced integration domain.
+Similarly, the effect of parameters $\beta_{0}$ and $\beta_{1}$ is overestimated to offset the mean bias induced by the reduced integration domain.
+However, $\xi_{2}$ is estimated more consistently because the true data-generating process assumes a faster temporal decay than the lag parameter used to generate the data.
+However, because estimates of the scale parameters are sporadic, stronger priors should be favoured if model inference and interpretability are of primary concern.
 With missing data, the numerical approximation introduces some bias; even with the true $(r,q)$, parameter estimates are positively biased and the bias is typically exacerbated under model misspecification.
-The top left panel in Figure 2 displays the distribution of the noise parameter $\sigma^{2}$.
-Misspecification of the mean process results in systematically higher estimates of $\sigma^{2}$.
-In the presence of missing data, estimated $\sigma^{2}$ is generally larger, reflecting increased uncertainty.
-Figure 2: Estimation of $\sigma^{2}$ and out-of-sample performance of simulated data measures using the $R^{2}$ metric, mean-squared error and coverage of the 95% prediction interval.
-Red boxplots indicate the results obtained using true radius and time lag values $(r=10,q=5)$, while blue $(\tilde{r}=15,\tilde{q}=8)$ and green $(\tilde{r}=5,\tilde{q}=2)$ illustrate the effect of model misspecification on predicitive performance.
-Despite these biases, the $R^{2}$ values remain relatively high, particularly when a reduced region of integration is selected.
-This pattern persists in the presence of missing data.
-Similarly, the mean-squared error is low relative to the performance under the true parameters, although misspecified models exhibit greater variability in out-of-sample performance.
-Under the full data setting, coverage of the reduced model is lower, with prediction interval widths ranging from 2.81 to 6.30, compared to the wider intervals of the extended model (2.75 to 8.44).
-Finally, even with missing data, out-of-sample performance remains strong, with coverage exceeding 95%, reflecting wider prediction intervals and higher uncertainty estimates.
+Figure 2: Out-of-sample performance of simulated data measured using the mean deviation of the predictions from the true values, and coverage of the 95% prediction interval.
+Red boxplots indicate the results obtained using true radius and time lag values $(r=10,q=5)$, while blue $(\tilde{r}=15,\tilde{q}=8)$ and green $(\tilde{r}=5,\tilde{q}=2)$ illustrate the effect of model misspecification on predictive performance.
+Under the full data setting, the out-of-sample deviations ($\hat{Y}(\mathbf{s},t)-Y(\mathbf{s},t)$) results in Figure 2 align with Corollary 1.
+When the kernel functions decay to zero within the true region of integration, selecting larger $\tilde{r}$ and $\tilde{q}$ reduces the mean bias compared to a smaller integration region.
+Additionally, predictions are underestimated when choosing smaller lags with lower coverage of the 95% prediction intervals.
+In the presence of missing data, we observe similar patterns in the deviations of the predictions from the true value.
+Most importantly, our proposed numerical approximation produces robust model predictions with high coverage of the prediction intervals.
+These results show that in practice choosing sufficiently large lag values ensures both reliable parameter estimation and predictions.
 4 Application to PM2.5 mapping in Northern California
 4.1 Data
-The dataset comprises fine particulate matter (PM2.5) measurements from 67 air quality monitoring stations across Northern California between 1st June 2020 and 31st October 2020, coinciding with the region’s peak wildfire season.
-Ground-based data were augmented with satellite data of aerosol optical depth (AOD) from the Multi-Angle Implementation of Atmospheric Correction (MAIAC) algorithm, which provides high-resolution (1̃ km x 1 km) gridded observations from the Moderate Resolution Imaging Spectroradiometer (MODIS) instruments aboard NASA’s Terra and Aqua satellites [Lyapustin et al., 2018].
-To account for wildfire smoke influence on PM2.5 concentrations, smoke plume data were obtained from the National Oceanic and Atmospheric Administration’s (NOAA) Hazard Mapping System (HMS), which delineates smoke plume extent based on satellite imagery [NOAA Satellite and Information Service, 2025].
-Together these data enabled a comprehensive characterization of spatiotemporal variation in PM2.5 and its association with wildfire smoke during the 2020 fire season in Northern California.
-Figure 3: Map of Northern California study region on two consecutive days in June 2020 showing: (a) PM2.5 monitoring locations (blue dots) with 20 km buffers around them (shaded blue circles) with 1 km x 1 km MAIAC aerosol optical depth (AOD) grid (pink to red); (b) zoom in of one PM2.5 monitor on a day with nearly completely observed AOD within 20 km; and (c) zoom in of one PM2.5 monitor on the next day with significant missing AOD within 20 km
+The dataset comprises of fine particulate matter (PM2.5) measurements from 67 air quality monitoring stations across Northern California between 1st June 2020 and 31st October 2020, coinciding with the region’s peak wildfire season.
+Ground-based data were linked with satellite data of aerosol optical depth (AOD) from the Multi-Angle Implementation of Atmospheric Correction (MAIAC) algorithm, which provides high-resolution (1̃ km x 1 km) gridded observations from an algorithm applied to observations from the Moderate Resolution Imaging Spectroradiometer (MODIS) instrument aboard NASA’s Terra and Aqua satellites (lyapustin2018).
+To account for wildfire smoke influence on PM2.5 concentrations, we included smoke plume data from the National Oceanic and Atmospheric Administration’s (NOAA) Hazard Mapping System (HMS), which delineates smoke plume extent based on satellite imagery (NOAA_HMS_Smoke).
+Together these data enabled the characterization of spatiotemporal variation in PM2.5 and its association with wildfire smoke during the 2020 fire season in Northern California.
+Figure 3: Map of Northern California study region on two consecutive days in June 2020 showing: (a) PM2.5 monitoring locations (blue dots) with 20 km buffers around them (shaded blue circles) with 1 km x 1 km MAIAC aerosol optical depth (AOD) grid (pink to red); (b) zoom in of one PM2.5 monitor on a day with nearly completely observed AOD within 20 km; and (c) zoom in of one PM2.5 monitor on the next day with significant missing AOD within 20 km.
 4.2 Model Specifications and Estimation
@@
 For the temporal kernel scale we used a weakly informative inverse-Gamma prior $\xi_{2}\sim\text{IG}(a_{0},b_{0})$.
-Building on the baseline model, the functional mean model augments the intercept by a temporally varying mean function $\beta_{0}(t)$ represented with B-spline basis functions $f_{k}(t)$ [Wood, 2017].
+Building on the baseline model, the functional mean model augments the intercept by a temporally varying mean function $\beta_{0}(t)$ represented with B-spline basis functions $f_{k}(t)$ (wood_generalized_2017).
 We used $K=10$ equally spaced knots to capture bi-weekly trends, and retained the same priors as in the baseline model.
@@
 Let $I(\mathbf{s},t)$ indicate a smoke plume event (1 = satellite-derived wildfire smoke plume intersected a PM2.5 monitor $Y(\mathbf{s},t)$ at location $s$ on day $t$, 0 = otherwise).
-We incorporated this indicator to the models: (1) at the mean level where $\beta_{0}(\mathbf{s},t)=\sum_{k}b_{jk}f_{k}(t)+I(\mathbf{s},t)$, (2) as a varying coefficient where the effect of the weighted predictor of AOD varied by plume status $\beta_{1}(\mathbf{s},t)=\beta_{10}+\beta_{11}I(\mathbf{s},t)$, and (3) by fixing $\xi_{2}$ at the posterior median from the functional mean model and allowing the spatial scale to depend on smoke plume status.
-This was motivated by [Zhang, 2004], who showed that range and scale parameters in covariance functions are not consistently estimable and require stabilization.
-The full model includes all features above and allows the spatial extent over which AOD influences PM2.5 to vary based on smoke plumes via $\log\{\xi_{1}(\mathbf{s},t)\}=\alpha_{0}+\alpha_{1}\,I(\mathbf{s},t)+\nu,\qquad\nu\sim N(0,\sigma_{\nu}^{2})$.
-To help estimation, when $I(\mathbf{s},t)=0$, we placed a strong point-mass prior on $\xi_{1}(\mathbf{s},t)$ centered at 2.0 for $r=10$ and 1.0 for $r=20$ (values obtained under the varying coefficient model).
+We incorporated this indicator to the models: (1) at the mean level where $\beta_{0}(\mathbf{s},t)=\sum_{k}b_{0k}f_{k}(t)+\gamma I(\mathbf{s},t)$ assuming a normal prior for $\gamma$, (2) as a varying coefficient where the effect of the weighted predictor of AOD varied by plume status $\beta_{1}(\mathbf{s},t)=\beta_{10}+\beta_{11}I(\mathbf{s},t)$.
+We fix $\xi_{2}$ at the posterior median from the functional mean model and allowing the effect of the kernel-averaged AOD to depend on smoke plume status.
+This was motivated by zhang_inconsistent_2004, who showed that range and scale parameters in covariance functions are not consistently estimable and require stabilization.
+The full model includes the varying coefficient specification above and allows the spatial extent over which AOD influences PM2.5 to vary based on smoke plumes via $\log\{\xi_{1}(\mathbf{s},t)\}=\alpha_{0}+\alpha_{1}\,I(\mathbf{s},t)+\nu,\qquad\nu\sim N(0,\sigma_{\nu}^{2})$.
+To help estimation, when $I(\mathbf{s},t)=0$, we placed a strong point-mass prior on $\xi_{1}(\mathbf{s},t)$ centered at the posterior median of the samples obtained under the varying coefficient model.
+We assume normal priors for $\mathbf{\alpha}$ and inverse-gamma for $\sigma_{\nu}^{2}$.
 Under this approach, we effectively only estimated the effect of $I(\mathbf{s},t)$ through $\alpha_{1}$.
-The MCMC algorithms used for inference can be found under Appendix D.
-Convergence was assessed using the Gelman-Rubin test and effective sample size.
-Each model was run for 25000 MCMC iterations, with the first 12500 samples used as warm-up and retaining the latter 12500 of the samples for inference.
+The MCMC algorithms used for inference can be found under Appendix D in the Supplementary Material.
+Convergence was assessed using the Gelman-Rubin test.
+Each model was run for 25000-50000 MCMC iterations, with the first 50% of the samples used as warm-up and retaining the latter half for inference.
 As a benchmark, we also fit a collocated linear model where we aggregated and averaged AOD data within a 5 km radius around each station on each day.
@@
 Varying coefficient $\sum_{k}b_{0k}\,f_{k}(t)$ $\gamma\,I(\mathbf{s},t)$ $\beta_{1}(\mathbf{s},t)=\beta_{10}$ $\beta_{11}I(\mathbf{s},t)$ $\xi_{1}\sim\mathrm{Gamma}(1,9)$ $\xi_{2}$ fixed at posterior median
-Full $\sum_{k}b_{0k}\,f_{k}(t)$ $\gamma\,I(\mathbf{s},t)$ $\beta_{1}(\mathbf{s},t)=\beta_{10}$ $\beta_{11}I(\mathbf{s},t)$ $\log\!\{\xi_{1}(\mathbf{s},t)\}=\alpha_{0}$ $\alpha_{1}I(\mathbf{s},t)+\nu,$ $\nu\sim N(0,\sigma_{\nu}^{2})$ $\xi_{2}$ fixed at posterior median
+Full $\sum_{k}b_{0k}\,f_{k}(t)$ $\beta_{1}(\mathbf{s},t)=\beta_{10}$ $\beta_{11}I(\mathbf{s},t)$ $\log\!\{\xi_{1}(\mathbf{s},t)\}=\alpha_{0}$ $\alpha_{1}I(\mathbf{s},t)+\nu,$ $\nu\sim N(0,\sigma_{\nu}^{2})$ $\xi_{2}$ fixed at posterior median
 Collocated Linear $\beta_{0}$ $\beta_{1}$
- Following [Zhang, 2004], $\xi_{2}$ is fixed at the posterior median from the functional mean model to stabilize range estimation.
+ Following zhang_inconsistent_2004, $\xi_{2}$ is fixed at the posterior median from the functional mean model to stabilize range estimation.
 In all models, $\beta_{1}(\mathbf{s},t)$ multiplies $W(\mathbf{s},t)$, the kernel-weighted AOD predictor defined in Eq.(1), which averages $X(\mathbf{u},t-l)$ over $\mathcal{B}_{r}(\mathbf{s})\times[0,q]$ using normalized exponential kernels in space and time.
-The collocated baseline model uses the AOD average $\overline{X}_{5\text{km}}(\mathbf{s},t)$ in 5 km buffer instead.
+The collocated linear benchmark model uses the AOD average $\overline{X}_{5\text{km}}(\mathbf{s},t)$ in 5 km buffer instead.
 We assumed independent Gaussian errors as our primary objective is to characterize the spatio-temporal dependence of PM2.5 on AOD.
@@
 4.3 Results
-The performance of each method is assessed using the root-mean square error (RMSE) metric, percentage of variance explained by the model (R2) and coverage of the 95% prediction interval for stations that were kept out-of-sample.
-The baseline collocated linear model produced an RMSE of 25.8, $R^{2}=0.184$ and coverage 0.941. To fit this model, 407 observations were removed due to unpaired PM2.5-AOD observations.
+The performance of each method is assessed using the root-mean square error (RMSE) metric, continuous ranked probability score (CRPS) to evaluate the accuracy of predictive forecasts (gneiting2007strictly) and the coverage of the 95% prediction interval for stations that were kept out-of-sample.
+The collocated linear model produced an RMSE of 25.8, CRPS 6.89 and coverage 0.941. To fit this model, 407 observations were removed due to unpaired PM2.5-AOD observations.
 Out-of-sample predictive performance of the competing models is reported in Table 2.
-Overall, the smaller buffer radius generally improved predictive accuracy, with the varying coefficient model benefitting less from this adjustment.
-In particular, the full model performs noticeably worse when the radius is set to $r=20$, suggesting that the added complexity at this radius may induce overfitting and reduce generalizability.
-However, we observed that the error is largely driven by overinflated predictions at a single location; when that location is removed, the model performance improves to RMSE = 19.3, R2 = 0.455, and coverage = 0.942.
-By contrast, when the radius was restricted to $r=10$, the full model achieved the best balance across evaluation metrics, outperforming the alternative specifications.
-These findings highlight the sensitivity of model performance to the choice of buffer radius and emphasize the importance of carefully selecting this parameter in practice.
+Overall, the larger radius around each site (r = 20 km) had better predictive accuracy (lower RMSE and CRPS), except for the full model, which performed better with r = 10 km.
+When $r=20$ the varying coefficient model achieved the best balance across evaluation metrics, outperforming the alternative model specifications.
 Table 2: Spatio-temporal model predictive performance for $r\in\{10,20\}$ km and $q=3$.
 r=10,q=3r=10,\;q=3 r=20,q=3r=20,\;q=3
-Model RMSE $\boldsymbol{R^{2}}$ Coverage RMSE $\boldsymbol{R^{2}}$ Coverage
-Baseline 32.5 0.230 0.956 35.1 0.227 0.957
-Functional mean 19.6 0.464 0.940 21.0 0.410 0.942
-Varying coefficient 21.0 0.362 0.944 19.9 0.400 0.944
-Full 19.0 0.454 0.940 63.3 0.037 0.940
-Collocated baseline model: RMSE = 25.8, $R^{2}=0.184$, Coverage = 0.941.
-To gain further insight into the spatio-temporal dependencies between PM2.5 and AOD, Figure 4 illustrates the estimated surface $\hat{\omega}(h,l)=\hat{\beta}_{1}\,\kappa_{1}(h;\hat{\xi}_{1}(\mathbf{s},t))\,\kappa_{2}(l;\hat{\xi}_{2}(\mathbf{s},t)),$ for the two best-performing models: the functional mean ST-Exp model (Figure 4(a)) and the full ST-Exp model (Figures 4(b) and 4(c)), both with $r=10$.
+Model RMSE CRPS Coverage RMSE CRPS Coverage
+Baseline 36.3 8.49 0.957 36.1 8.22 0.953
+Functional mean 19.6 6.32 0.942 20.1 6.21 0.945
+Varying coefficient 23.5 6.97 0.942 19.1 6.07 0.946
+Full 19.6 6.22 0.943 20.1 6.34 0.944
+[…]
+To gain further insight into the spatio-temporal dependencies between PM2.5 and AOD, Figure 4 illustrates the estimated surface $\hat{\omega}(h,l)=\hat{\beta}_{1}\,\kappa_{1}(h;\hat{\xi}_{1}(\mathbf{s},t))\,\kappa_{2}(l;\hat{\xi}_{2}(\mathbf{s},t)),$ for the three best-performing models: the functional mean model (Figure 4(a)), the varying coefficient model (Figures 4(b) and 4(c)), both with $r=20$, and the full model (Figures 4(d) and 4(e)) with $r=10$.
 We used the posterior medians of the MCMC samples to construct the surface $\widehat{\omega}(h,l)$.
-Figure 4(b) shows the estimated effect of AOD on PM2.5 when no smoke plume was present ($I(\mathbf{s},t)=0$), whereas Figure 4(c) shows the effect under plume conditions ($I(\mathbf{s},t)=1$).
-The results indicate that incorporating wildfire smoke information alters the spatio-temporal weighting: in the functional mean model, the AOD effect was distributed more evenly across the buffer region, while in the full model, both for $I(\mathbf{s},t)=0$ and $I(\mathbf{s},t)=1$, the effect decayed more rapidly over space and time but attained larger magnitudes in the immediate neighbourhood of an observation.
-Taken together, these results suggest that fire events amplify the local influence of AOD on PM2.5, underscoring the importance of accounting for such events when constructing accurate exposure maps.
+Figures 4(b) and  4(d) show the estimated effect of AOD on PM2.5 when no smoke plume was present ($I(\mathbf{s},t)=0$), whereas Figure 4(c) and  4(e) show the effect under plume conditions ($I(\mathbf{s},t)=1$).
+The results indicate that incorporating wildfire smoke information alters the spatio-temporal weighting.
+The varying coefficient model is able to isolate the effect of AOD under different plume conditions.
+In the presence of fire plumes, the model estimates higher levels of PM2.5 with $\hat{\gamma}\in[0.461,0.587]$ on average while the effect of the kernel-averaged AOD being smaller and more localised in its immediate neighbourhood.
+The full model captures this differently, the magnitude of the effect of the weighted AOD on PM2.5 is higher in the presence of smoke plumes.
+These results suggest that fire events amplify the local influence of AOD on PM2.5, underscoring the importance of accounting for such events when constructing accurate exposure maps.
 (a) Functional mean
-(b) Full model in the absence of fire plumes
-(c) Full model in the presence of fire plumes
-Figure 4: Estimates of the surface explaining the spatio-temporal dependencies between PM2.5 and AOD at varying spatial and time lags ($r=10$).
+($r=20$)
+(b) Varying coefficient model in the
+absence of fire plumes ($r=20$)
+(c) Varying coefficient model in the presence of fire plumes ($r=20$)
+(d) Full model in the absence of
+fire plumes ($r=10$)
+(e) Full model in the presence of fire plumes ($r=10$)
+Figure 4: Estimates of the surface explaining the spatio-temporal dependencies between PM2.5 and AOD at varying spatial and time lags.
 Figure 5: Predictions of PM2.5 at out-of-sample locations over time.
-The top panel compares the collocated linear model with the functional mean model ($r=10$).
-The bottom panel compares the functional mean model with the full model ($r=10$).
+The top panel compares the collocated linear model with the baseline model ($r=20$).
+The bottom panel compares the baseline model with the varying coefficient model ($r=20$).
 As model complexity increases, uncertainty quantification of the predictions improves with narrower prediction intervals capturing peaks.
@@
 The collocated linear benchmark model attained reasonable point predictions but produced very wide intervals, especially during sharp concentration peaks.
-By contrast, the functional mean model has substantially narrower interval widths while maintaining coverage, providing more precise uncertainty quantification for PM2.5 concentrations across both peak episodes and stable periods.
+By contrast, the baseline model has substantially narrower interval widths while maintaining coverage, providing more precise uncertainty quantification for PM2.5 concentrations across both peak episodes and stable periods.
 Figure 5(b) further shows that incorporating wildfire information improves temporal calibration.
 The full model produced narrower prediction intervals that still captured the observed peaks, highlighting the importance for wildfire influences in PM2.5 estimation.
-Figure 6 presents 5 km resolution gridded maps of estimated PM2.5 derived from 1 km resolution AOD for three days in September during peak wildfire season, using the functional mean model.
+Figure 6 presents 1 km resolution gridded maps of estimated PM2.5 derived from 1 km resolution AOD for three days in September during peak wildfire season, using the functional mean model.
 The approach effectively captured both spatial gradients and day-to-day temporal patterns, producing smooth, coherent variation across space and time, particularly where the August Complex wildfire occurred.
-Figure 6: Spatial mapping of PM2.5 using the functional mean model for three consecutive days in September 2020.
+An advantage of our approach is it enables fine-resolution predictions even when AOD is missing.
+Figure 6: Spatial mapping of PM2.5 using the full model ($r=10$) for three consecutive days in September 2020.
 5 Discussion
-Handling spatial misalignment traditionally relies on change-of-support or hierarchical modeling to reconcile point- and areal-level data [Gotway and Young, 2002].
+Handling spatial misalignment traditionally relies on change-of-support or hierarchical modeling to reconcile point- and areal-level data (gotway_young_2002).
 In point-to-point and point-to-grid misalignment, common practice is to average data from the misaligned source over buffer regions around a target point to align (i.e.
@@
 Unlike common strategies that first gap-fill (interpolate) and then regress on the complete interpolated surface (e.g.
-[Li and Heap, 2014]), our approach builds the weighting and uncertainty handling directly into the likelihood via $W(\mathbf{s},t)$, allowing propagation of uncertainty from the available data to the predictions.
+li2014), our approach builds the weighting and uncertainty handling directly into the likelihood via $W(\mathbf{s},t)$, allowing propagation of uncertainty from the available data to the predictions.
 Missingness and and spatial misalignment are common issues in many environmental modeling applications.
-In the satellite AOD–PM2.5 literature, missing AOD due to clouds, snow, high aerosol loadings, or retrieval failures is widespread, and many studies use explicit gap filling via interpolation or ensembles before regression [Xiao et al., 2021, Sorek-Hamer et al., 2020, Li et al., 2020].
-When we applied our approach to estimate PM2.5 from satellite AOD over Northern California in 2020, which had an intense wildfire season, the weighted predictor model captured both spatial and temporal variation and yielded narrower, better-calibrated prediction intervals than a collocated linear baseline, with more than a 50% increase in R2.
+In the satellite AOD–PM2.5 literature, missing AOD due to clouds, snow, high aerosol loadings, or retrieval failures is widespread, and many studies use explicit gap filling via interpolation or ensembles before regression (xiao2021; sorek-hamer2020; li2020).
+When we applied our approach to estimate PM2.5 from satellite AOD over Northern California in 2020, which had an intense wildfire season, the weighted predictor model captured both spatial and temporal variation and yielded narrower, better-calibrated prediction intervals than a collocated linear baseline.
 Incorporating smoke plume information altered the estimated weighting surface and improved temporal calibration around concentration peaks.
@@
 Future work will integrate multi-source covariates (meteorology, emissions, chemical transport model outputs) and explore adaptive, plume-aware kernels to further enhance predictive performance and interpretability.
+[…]
 Appendix A Voronoi Tessellation
+Figure  7 shows the Voronoi tessellation over a buffer of 20km for different days. On days where there is missing AOD data, the region is weighted more heavily due to a larger Voronoi cell.
+Appendix B Simulation Study
 Figure 7: Example of a Voronoi tessellation over a buffer.
-Appendix B Simulation Study
-Figure 8: Examples of simulated $X(s,t)$ fields.
 Figure 8: Examples of weight surface used in simulation study.
@@
     Thus
-    $\bigl|\mu(\mathbf{s},t)-\tilde{\mu}(\mathbf{s},t)\bigr|\leq\beta\int_{0}^{q}\int_{\mathcal{A}_{r,\tilde{r}}(\mathbf{s})}\kappa_{1}(\mathbf{s}-\mathbf{u})\,\kappa_{2}(t-\ell)\,X(\mathbf{u},t-\ell)\,d\mathbf{u}\,d\ell.$
+    $\bigl|\mu(\mathbf{s},t)-\tilde{\mu}(\mathbf{s},t)\bigr|\leq\beta_{1}\int_{0}^{q}\int_{\mathcal{A}_{r,\tilde{r}}(\mathbf{s})}\kappa_{1}(\mathbf{s}-\mathbf{u})\,\kappa_{2}(t-\ell)\,X(\mathbf{u},t-\ell)\,d\mathbf{u}\,d\ell.$
   2. (2)
@@
     over the common ball $\mathcal{B}_{r}(\mathbf{s})$. Hence
-    $\bigl|\mu(\mathbf{s},t)-\tilde{\mu}(\mathbf{s},t)\bigr|\leq\beta\int_{\mathcal{B}_{r}(\mathbf{s})}\int_{\min(q,\tilde{q})}^{\max(q,\tilde{q})}\kappa_{1}(\mathbf{s}-\mathbf{u})\,\kappa_{2}(t-\ell)\,X(\mathbf{u},t-\ell)\,d\ell\,d\mathbf{u}.$
+    $\bigl|\mu(\mathbf{s},t)-\tilde{\mu}(\mathbf{s},t)\bigr|\leq\beta_{1}\int_{\mathcal{B}_{r}(\mathbf{s})}\int_{\min(q,\tilde{q})}^{\max(q,\tilde{q})}\kappa_{1}(\mathbf{s}-\mathbf{u})\,\kappa_{2}(t-\ell)\,X(\mathbf{u},t-\ell)\,d\ell\,d\mathbf{u}.$
   3. (3)
@@
     $A_{2}=\mathcal{A}_{r,\tilde{r}}(\mathbf{s})\times[0,\,\tilde{q}$].
-    Thus the absolute difference of the two integrals equals the integral of the (nonnegative) integrand over $A_{2}$ minus the integral over $A_{1}$ in magnitude, and is therefore bounded above by the integral of the integrand over
+    Thus the absolute difference of the two integrals equals the integral of the (nonnegative) integrand over $A_{2}$ minus the integral over $A_{1}$ in magnitude, and is therefore bounded above by the integral over
     $A_{1}\cup A_{2}.$
@@
 Appendix D Brief description of MCMC algorithm
-Under a Gaussian likelihood, the conditional posterior distribution $\beta|\cdot$ is given by
+Let $\beta$ denote the regression coefficients $\mathbf{b},\gamma]$ Under a Gaussian likelihood, the conditional posterior distribution $\beta|\cdot$ is given by
 $\mathbf{\beta}|\mathbf{\xi},\sigma^{2}\propto\pi(\mathbf{\beta})\times\prod_{i,j}f(Y(s_{i},t_{j})\;;W(s_{i},t_{j}\;;\xi(s_{i},t_{j})),\sigma^{2})$ (A1.1)
@@
 end for
-Algorithm 2 Metropolis-within-Gibbs when $\log(\xi_{m}(\mathbf{s},t))=\alpha_{0}+\ \alpha_{1}I(s,t)+\nu$
+Algorithm 2 Metropolis-within-Gibbs when $\log(\xi_{1}(\mathbf{s},t))=\alpha_{0}+\ \alpha_{1}I(s,t)+\nu$